# An Optimal, Universal and Agnostic Decoding Method for Message Reconstruction, Bio and Technosignature Detection

### v3

@@ Abstract
We present a signal reconstruction method for zero-knowledge one-way communication channels in which a receiver aims to interpret a message sent by an unknown source about which no prior knowledge is available and to which no return message can be sent. Our reconstruction method is agnostic vis-à-vis the arbitrarily chosen encoding-decoding scheme and other observer-dependent characteristics, such as the arbitrarily chosen computation model or underlying mathematical theory. We investigate how non-random messages may encode information about the physical properties, such as dimension and length scales of the space in which a signal or message may have been originally encoded, embedded, or generated. We argue that our results have applications to life and technosignature detection and to coding theory in general.
@@
algorithmic information dynamics, causal deconvolution, zero-knowledge communication, technosignatures, signal processing, perturbation analysis, universal distribution, intelligent signal detection, biosignatures
@@ 1 Introduction
For the past 50 years, astronomers have been sweeping the skies using radio telescopes in hopes of stumbling across a message from an alien civilisation. More recently, space missions to collect samples and data from Mars, Titan, Europa, and others have been focused on detecting possible biosignatures [1, 2, 3, 4, 5]. The James Webb Space Telescope has been outfitted with sensors and measurement devices to scout potentially habitable solar systems [6, 7]. For the first time in human history, the detection of extraterrestrial life is at the forefront of space exploration. In the novel Contact, written by Carl Sagan and later made into a film, an extraterrestrial signal is received that is encoded in three dimensions. A good deal of the story has to do with how scientists serendipitously deduce that the signal encodes an object in three dimensions after spending months or years trying to fit it in two dimensions.
Here we show that the more a message is removed from randomness the easier is to derive its native geometric and topological dimensions, and that such an approach derives its fruitful properties from the universality and limitations of algorithmic compression algorithms. These properties include being agnostic (in the asymptotic limit when computational resources are unbounded) vis-á-vis: the encoding-decoding scheme used to encode the message by the emitter agent; computation model and the programming language chosen in order to implement our proposed method; the computable (or semi-computable) method of approximation to algorithmic complexity; any arbitrarily chosen (computable) probability measure of the events; and the formal theory able to formalize the mathematical statements. While the arbitrary choice of each of these constitutes an observer dependency, this dependency is cancelled out in the long run without the need of resorting to special observers [8].
@@
All of our knowledge about life is based on what occurs here on Earth [9, 10, 11, 12, 13, 14, 15, 16, 17]. Some researchers consider technosignature detection a more promising method of detecting life than other biosignatures because of their longevity [18, 19, 20]. However, some researchers worry that our current theoretical toolkit for data analysis may not be sophisticated enough [21, 22, 23]. Most of the discussion has been centred around the technical justification of hardware choices and the technicalities of detection (e.g. the frequency) and not the nature (i.e., original structural characteristics or underlying mathematical properties) of the signal itself. Thus, for example, the search for extraterrestrial signals has been mostly focused on narrow-band signals (mostly pulses) a few Hertz-wide (or narrower). Natural cosmic noisemakers, such as pulsars, quasars, and the turbulent, thin interstellar gas of our own Milky Way do not make radio signals that are this narrow, which is a reasonable justification of the bounds of the physical signal search. On Earth, we also find local signal-noise makers such as animals, organs, neurons, and cells. However, beyond identifying very basic statistical patterns, not much progress has been made on the qualitative semantic aspects of signals and communication, let alone on building a universal framework within which the question can be explored and analysed.
@@
Each different species on Earth has its own different set of sensors that it uses to interact with its environment. To put this in terms of signals or messages between senders and receivers, different species emit signals based on the mechanisms that allow them to do so [24, 25]. They also decode messages based on their sensory capabilities, and if each species has a unique set of senses, then we can say that each species decodes the same message differently. The 52Hz whale’s messages are decoded differently between humans monitoring the oceans and other whales (who do not sense the messages at all).
@@
One can say that that if a signal is intentional, it might be decipherable. In order to intentionally send or receive a signal over interstellar distances, it is reasonable to assume a civilisation must understand basic science and mathematics, or at least must evolve in such that these are somehow encoded into a message. Hence, a message from another civilisation might use some similar framework to science and mathematics to build up a common ground with respect to other societies. Signals sent by a civilisation for its own purposes may be impossible for us to unravel but our own mathematical tools may help.
@@
In this paper, we introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (i.e., multidimensional space) in which the original message was given its “meaning” by the emitter agent. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are the objects grounded in their respective contexts, that is, embedded into their respective multidimensional spaces as the emitter originally intended. See the Sup. Inf. for a formal introduction to these concepts.
@@
One-way communication channels are those for which the receiver cannot (in principle or in practice) send any signal back to the emitter in order to help or facilitate the decoding process of the first message sent by the emitter. We demonstrate how signals and messages may be reconstructed by deriving the number of dimensions and the scale of each length of an object from examples ranging from text to images embedded in multiple dimensions, showing a connection between irreducible information content, syntax, geometry, semantics, and topology. In addition, this article also shows results that are not only agnostic vis-á-vis prior knowledge of encoding-decoding schemes, but also demonstrate sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
@@
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes, and consequentially also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Sup. Inf., this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [26, 27, 28]. Actually, one may consider ZKP and ZKC as kindred mathematical problems but as diametrically opposing counterparts with regard to the acquisition of knowledge.
@@
The method is based on the principles of Algorithmic Information Dynamics (AID) [29, 30, 8], and it consists of a perturbation analysis of a received signal. AID is based upon the principles of information theory and the mechanisms of algorithmic probability and the universal distribution, a formal approach to a type of Artificial General/Super Intelligence that requires the massive production of a universal distribution, the mother of all models [31], to build a very large (semi-computable) model of computable approximate models. The underlying idea is that a computable model is a causal explanation of a piece of data. As a result of each perturbation applied to the received signal, a new computable approximate model is built, and then compared against the observation [30, 32]. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. In the particular context investigated in this article, the partition (i.e., multidimensional space) for which the message displays the lowest complexity indicates the original partition that the received signal stream encodes.
@@
Overarching frameworks such as the one introduced here can be useful not only for signal detection but also for signal deconvolution from local information in, e.g., biology (what chemical signals among cells mean). In previous work, we showed how the same technology can be used to disentangle the 3D structure of DNA and genomic information [33].
2 Perturbation analysis of compressed data
To showcase the relationship between compression, algorithmic complexity, and encoding methods, introducing the method put forward in this work, we took real sentences from Darwin’s “Origin of Species” and binarised them in two different ways according to arbitrary ASCII-to-binary functions. In one encoding (left column in Fig. 1), ASCII vowels are assigned a 1 and all others 0. For the other encoding (right column), spaces are assigned a 1 and the rest 0. The compression, block entropy, and block decomposition method (BDM) are measured for each string. Then, for each string, each value is perturbed (1 to a 0, or vice-versa) one at a time. The resulting change in values from the original binary sequence to the perturbed ones are shown as different coloured lines for each panel.
The results in Fig. 1 demonstrate that most single-bit perturbations in a sequence that encodes a message result in a more random-looking sequence. These results vary depending on what mapping we choose to binarise the sequences, but in general, any perturbation generally increases randomness across various mappings (we also explored other arbitrary binarisation functions, with similar results). In other words, this example shows that sequences that carry meaningful content are quantitatively different according to various measures. Any deviation (i.e., perturbation) from the encoded sequences causes the randomness to increase, which suggests that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
Fig. 2 demonstrates the same phenomenon shown in non-random text but in audio signals, the results using an audio recording of the words spoken from Apollo 13, “Houston, we’ve had a problem” transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. When compressed, the message is 106 250 bytes and 106 320 characters long. Here, the message was scrambled several times. The histograms on the bottom left show the different compression lengths in bytes and string lengths of the scrambled messages. Each of these scramblings results in an increase in randomness. Additionally, if only the beginning and end of each word are perturbed, the resulting randomness increases smoothly with each perturbation, as shown in the bottom right plot. These results indicate that perturbations of the original message, including scrambling elements of the original message in a different order, will result in a message of higher complexity and thus greater randomness. As a result, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity.
@@ 3 On the information content of the Arecibo message
In 1974, the bitmap image in Fig. 3 was sent into space as a radio signal from the Arecibo radio telescope. At the left-hand end of the image is a version of the pattern of digits from page 117– but it is distorted so it has no obvious nested structure. In the image, there are atomic numbers for various elements and bitvectors for components of DNA. Under these, there are rough pictorial representations of a DNA molecule, a human, and the telescope. All these images seem to depend almost completely on human visual encoding/decoding conventions. Without any sort of human context, including any indication that these are pictorial representations based on human vision, their meaning would be essentially impossible to recognise. This is especially true for message receivers who may not possess visual recognition capabilities, at least not visual capabilities that are similar to our own vision.
@@
In addition, this message was reformatted from its original bidimensional state and sent as a string of 1,679 binary digits. The reasoning was that an alien civilisation with a mathematical system that receives the message will recognise 1,679 as a semi-prime number–a multiple of 23 and 73. Even if the original message could not be reconstructed in its original and intended encoding, the receiver would recognise the message as having some mathematical significance that is unlikely to be a random signal from outer space.
@@
Fig. 4 shows how the method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space, by sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [34] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible, and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
@@
Fig. 5 shows the method’s noise resistance to indicate the candidate partition and the precise lengths of the original message. Successful amplification of the signal is shown in Fig. 6, with BDM outperforming Compression and Shannon Entropy in the face of additive noise, with Compression showing insensitivity to the original signal at about 10% of bits flipped, and Shannon Entropy diminishing faster than BDM but slower than Compression. In all cases, different levels of robustness at deconvolving the image dimensions are shown.
@@
Investigating this further, Fig. 7 shows six bidimensional images along with their original numerical dimension. The values of the three complexity measures over possible partitions are shown below the original numerical dimension for ease of comparison. The method is invariant with regard to linear transformations such as encoding, as shown in images 2 and 4 of Fig. 7. Drops in complexity (downward spikes) indicate candidate dimensions for the original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. BDM outperforms compression and entropy and correctly identifies the original message (image), encoding for 4 out of 6 images.
@@
Some of these image decodings for the space shuttle image are shown in Fig. 8. Each panel shows random realignments of the correct dimensions along with the original image (i.e., object) embedded into these new partitions, with their respective BDM values displayed above each image. When compared to the other bidimensional spaces, the one that is most similar to the original message will have a lower complexity value.

### v4

@@ Abstract
We present an agnostic signal reconstruction method for zero-knowledge one-way communication channels in which a receiver aims to interpret a message sent by an unknown source about which no prior knowledge is available and to which no return message can be sent. Our reconstruction method is agnostic vis-à-vis the arbitrarily chosen encoding-decoding scheme and other observer-dependent characteristics, such as the arbitrarily chosen computational model, probability distributions, or underlying mathematical theory. We investigate how non-random messages encode information about their intended physical properties, such as dimension and length scales of the space in which a signal or message may have been originally encoded, embedded, or generated. We focus on image data as a first illustration of the capabilities of the new method. We argue that our results have applications to life and technosignature detection, and to coding theory in general.
@@
algorithmic information dynamics, causal deconvolution, zero-knowledge communication, technosignatures, signal processing, perturbation analysis, universal distribution, intelligent signal detection, meaning, semantics, geometry, biosignatures
@@ 1 Introduction
The interpretation of any given message may be thought of as highly observer-dependent. This is because the meaning of a message can vary from one recipient to another, and indeed over time. However, not all signals are the same. For example, signals with fewer symbols can be combined in fewer possible messages implying they are potentially restricted in the semantic space of possible meanings. Random signals with large size vocabularies may have also more ways to be combined to cover a larger semantic space. If originally intended to have meaning by the emitter agent (or sender), truly random signals will require a lot of effort—i.e., a sufficiently high amount of computational resources to compute/decode the respective algorithmic information content expressed in its incompressible form—to produce structured messages in any semantic context at the other end of the receiver agent, otherwise they would not be (algorithmic) random.
@@
However, even signals such as the cosmic microwave background from all directions in the universe may later amount for the greatest scientific evidence that our universe had a beginning, as an example of possible white noise converted into a thermal signature of highly contextual meaning. For the past 50 years astronomers have been sweeping the skies using radio telescopes in search for signals of possible extraterrestrial intelligence origins that would be meaningful to us with no reproducible success. More recently, space missions collecting samples and data from Mars, Titan, Europa, and other extraterrestrial bodies have been focused on detecting possible biosignatures with interesting results. And some tools have been developed with different degrees of relevancy and success [1, 2, 3, 4, 5, 6]. The James Webb Space Telescope, for example, has been fitted with sensors and measurement devices to scout potentially habitable solar systems [7, 8]. For the first time in human history, the detection of extraterrestrial biosignatures and life is at the forefront of space exploration. In this work, we ask whether some properties of messages, in special images, may be more objective even when they may have been arbitrarily constructed by their senders. In the novel Contact, written by Carl Sagan and later made into a film, an extraterrestrial signal was received. A good deal of the story has to do with how scientists serendipitously deduce that the signal encodes an object in three dimensions, having spent months or years trying to fit it into two dimensions.
@@
All of our knowledge about life is based on what occurs here on Earth [9, 10, 11, 12, 13, 14, 15, 16, 17]. Some researchers consider technosignature detection a more promising method of detecting life than other biosignatures because of technosignatures’ putative longevity [18, 19, 20]. However, some researchers worry that our current theoretical toolkit for data analysis may not be sophisticated enough [21, 22, 23]. Most of the discussion thus far has been centred around the technical justification of hardware choices and the technicalities of detection (e.g. the frequency) and not on the nature (i.e., original structural characteristics or underlying mathematical properties) of the signal itself. Thus, for example, the search for extraterrestrial signals has been mostly focused on narrow-band signals (mostly pulses) a few Hertz wide (or narrower). Natural cosmic noisemakers, such as pulsars, quasars, and the turbulent, thin interstellar gas of our own Milky Way do not make radio signals that are this narrow, which would be a reasonable justification of the bounds of the physical signal search. On Earth, we also find local signal-noise makers such as animals, organs, neurons, and cells. However, beyond identifying very basic statistical patterns, not much progress has been made on the qualitative semantic aspects of signals and communication, let alone on building a universal framework within which the question can be explored and analysed.
@@
Each different species on Earth has its own different set of sensors that it uses to interact with its environment. To put this in terms of signals or messages between senders and receivers, different species emit signals based on the mechanisms that allow them to do so [24, 25]. They also decode messages based on their sensory capabilities, and if each species has a unique set of senses, then we can say that each species decodes the same message differently. The 52Hz whale’s messages are decoded differently by humans monitoring the oceans and by other whales (who do not sense the messages at all).
@@
One can say that if a signal is intentional, it might be decipherable. In order to intentionally send or receive a signal over interstellar distances, it is reasonable to assume a civilisation must understand basic science and mathematics, or at least must evolve in such a way that these are somehow encoded into a message. Hence, a message from another civilisation might use a framework similar to human-based science and mathematics to establish common ground with respect to other societies. Signals sent by a civilisation for its own purposes other than to also establish that “common ground” may be impossible for a third party to unravel, though our own mathematical tools may help.
@@
In this paper, we introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (i.e., multidimensional space) in which the original message was given its “meaning” by the emitter agent (i.e., the sender intentionally aiming at communicating). Thus, we establish—for the first time, to our knowledge—a strong and deep connection between semantics, geometry and dimensional topology. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are images grounded in their respective contexts, that is, embedded into their respective multidimensional spaces as the emitter originally intended. See the Sup. Inf. and [26] for a formal introduction to these concepts.
@@
One-way communication channels are those for which the receiver cannot (in principle or in practice) send any signal back to the emitter in order to help or facilitate the decoding process of the first message sent by the emitter. We demonstrate how signals and messages may be reconstructed by deriving the number of dimensions and the scale of each length of an object from examples of images embedded in multiple dimensions, showing a connection between irreducible information content, geometry, topology, syntax and semantics. In addition, this article also presents results that are not only agnostic vis-á-vis prior knowledge of encoding-decoding schemes, but also demonstrate sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
@@
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of ’no prior information about the emitter agent’ only applies to encoding-decoding schemes, computation models, programming languages, probability measures, underlying formal mathematical theory, and also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Sup. Inf. and [26], this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. For example, the assumption that the emitter agent is somehow capable of performing an arbitrary encoding and compression of the message into a signal stream; and the assumption that the sender (or emitter agent) is not trying—e.g., by employing all of its computational power to outpace that of the receiver—to deceive the receiver agent into the wrong message, as both ends of the communication channel should at least have the intention to reach the above mentioned “common ground” although completely independent of each other. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [27, 28, 29]. Actually, one may consider ZKP and ZKC as kindred mathematical problems but as diametrically opposed counterparts with regard to the acquisition of knowledge [26].
@@
As pointed out, our method is based on the principles of Algorithmic Information Dynamics (AID) [30, 31, 32], and it consists of a perturbation analysis of a received signal. AID is based upon the principles of information theory and the mechanisms of algorithmic probability and the universal distribution, a formal approach to a type of Artificial General/Super Intelligence that requires the massive production of a universal distribution, the mother of all models [33], to build a very large (semi-computable) model of computable approximate models. The underlying idea is that a computable model is a causal explanation of a piece of data. As a result of each perturbation applied to the received signal, a new computable approximate model is built, and then compared against the observation [31, 34]. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. In the particular context investigated in this article, the partition (i.e., multidimensional space) for which the message (image) displays sufficiently low complexity values are candidate dimensions to match the original partition that the received signal stream encodes.
@@
Overarching frameworks such as the one introduced here can be useful not only for signal detection but also for signal deconvolution from local information in, e.g., biology (what chemical signals among cells mean). In previous work, we showed how the same technology can be used to disentangle the 3D structure of DNA and genomic information [35]. As introduced and formally demonstrated in [26], we show that it is an observer limitation to open-endedly approaching the theory that counter-intuitively allows the correct decoding of the sender’s original encoding, but that even when subjective, the process is universal, asymptotically agnostic, mathematically optimal, and independent of computational model, programming language, or probability distributions (including either priors or updated distributions).
@@ 2 On the information content of the Arecibo message
In 1974, the bitmap image on the left-hand side of Fig. 1 was sent into space as a radio signal from the Arecibo radio telescope. In the image, there are atomic numbers for various elements and bitvectors for components of DNA. Under these, there are rough pictorial representations of a DNA molecule, a human, and the telescope. All these images seem to depend almost completely on human visual encoding/decoding conventions. On the right-hand side of Fig. 1 is a reshaped version of the pattern of digits, but it is distorted so it has no obvious nested structure. Without any sort of human context, including any indication that these are pictorial representations based on human vision, their meaning would be essentially impossible to recognise. This is especially true for message receivers who may not possess visual recognition capabilities, at least not visual capabilities that are similar to our own.
Refer to caption           Refer to caption
@@
This message was reformatted from its original bidimensional state and sent as a string of 1,679 binary digits. The reasoning was that an alien civilisation with a mathematical system that receives the message will recognise 1,679 as a semi-prime number—a multiple of 23 and 73. Even if the original message could not be reconstructed in its original and intended encoding, the receiver would recognise the message as having some mathematical significance that is unlikely to be a random signal from outer space. In order to assess the information content change after structural perturbations, three metrics are measured for each perturbed array:
  • Entropy: Shannon entropy value of the binary array (which could be Block Entropy for higher dimensional data);
  • BDM: The Block Decomposition Method (BDM) [36] is a strategy used to approximate the algorithmic or Kolmogorov complexity of objects like strings and graphs. Unlike conventional methods that focus on compressibility, such as lossless compression techniques, BDM divides a complex object into smaller, easier-to-handle components. These individual blocks are then evaluated based on their complexity using a reference database of small Turing machine outputs. This database provides a baseline for assessing the algorithmic content of each block. BDM builds upon the Coding Theorem Method (CTM), which relates algorithmic probability to Kolmogorov complexity by determining the chance that a random program could produce a given string on a universal Turing machine. By applying this logic to each block and summing their complexities, BDM estimates the complexity of the whole object. This approach is especially practical when calculating exact Kolmogorov complexity is too difficult due to the object’s size. By combining the complexity estimates of smaller blocks, BDM offers a broader, more refined view of complexity, surpassing the limitations of basic compressibility measures like Shannon entropy. This method reveals deeper insights into both structured and random data patterns.
  • z-lib: the length of the b64encode of the compressed zlib version of the binary array (as a string). This is an alternative to LZW lossless compression.
The perturbations consist of rearranging the 1D stream as a 2D $m\times n$ array. Since the size of the original binary message $s=1679$ may not necessarily be reshaped as $m\times n$ (i.e., the original size is not the product of $m$ and $n$), it is important to set the maximum information loss allowed in this reshaping process. This can be set, for example as a 1 % loss for our investigation purposes, which indicates that we are looking for $m$ and $n$ such that $mn\geq 0.99s$, where $s$ is the length of the original encoded binary message. Also, there may be more than one pair of $m$ and $n$ values which satisfy such inequality, thus we set that $n$ is the maximum second dimension for a given $m$ value. So the dimensions actually searched are $(m,n)$, both natural numbers, such that $n=max(n_{1},n_{2},...n_{j})$, where $mn_{i}\geq 0.99s$, for $i=1,...,j$.
The following changes are then considered for each of the metrics chosen to be evaluated:
  • the BDM value is divided by the number of 4x4 grids used to make the partitioning of the original 2D array (i.e. $(m-3)(n-3)$);
  • Entropy value is changed to Block Entropy, where the block size is equal to $n$;
  • zlib b64encoding length is divided by the ratio of bits kept (i.e. $mn/s$).
To make the comparisons more clear, for every metric, a MinMaxScaler was applied, such that the results range from 0 to 1. Fig. 2 presents the (structural) perturbations [26] on the original Arecibo message 1D stream of binary digits.
Figure 2: Perturbations on the original 1D stream of binary digits of the Arecibo message. The “correct” shape of the message is represented in red and is observed for downward BDM peaks in the low-BDM region.
In general, BDM is still showing a high sensitivity to the change in information content. Block Entropy also presented a peak around the “correct” dimension. For small strings, z-lib may not perform well due to the overhead needed to decompress the strings in a lossless manner. Therefore, it may be of interest to upscale small strings. This could be done, for example, by changing a single bit to a square array of bits of the same value.
@@
Fig. 3 graphically shows how the method is robust when it comes to identifying the best way to decode a signal back to its original bidimensional space by sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages. After this process, the candidate dimension would be among the lowest algorithmic complexity configurations, under the assumption that the original message is not algorithmically random (i.e., that the original message is compressible). As shown in [37] for graphs, this happens because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very small. On the other hand, if the original message is sufficiently compressible, and therefore not random, most partitions will lead to configurations of greater algorithmic randomness. See [26] for the formal proofs in the general case.
Figure 3: Top left: Most possible partitions result in random-appearing configurations with high corresponding complexity, indicating measurable randomness. Bottom: Some partitions ($n$ values) will approximate the originally encoded meaning (third from the right). Other configurations result in images with higher complexity values. This sequence of images shows the images in the approximate vicinity of the correct bidimensional configuration (i.e., partition) and illustrates fast convergence to low complexity. Top right: By using different information indexes across different configurations, a downward-pointing spike will indicate message (image) configurations that correspond to low-complexity image(s). In this plot, log complexity index is just a version of a scaled version of each metric, created to provide better visibility of the spikes. This allows a prior-knowledge-agnostic and objective method to infer a message’s original encoding. Of the various measures, BDM, combining classical information (entropy) for long ranges and a measure motivated by algorithmic probability for short ranges, is the most sensitive and accurate in this regard. Traditional compression and entropy may also contribute to finding the right configuration amongst the top spiking candidates. The ratio of noise-to-signal was amplified in favour of the hidden structure by multiplying the original image size by 6 for both length and height (such amplification was necessary to make sure compression algorithms do not bias their results from the overhead needed to decode the messages).
@@
Fig. 4 shows the method’s noise resistance to indicate the candidate partition and the precise lengths of the original message. For this plot, normalised log complexity index is a scaled version of the metrics considered. Signal deconvolution before and after amplification is shown in Fig. 5, with BDM outperforming Compression and Shannon Entropy in the face of additive noise, with Compression showing insensitivity to the original signal at about 10% of bits flipped, and Shannon Entropy diminishing faster than BDM but slower than Compression. In all cases, different levels of robustness at deconvolving the image dimensions are evinced.
Figure 5: After applying 3 different quantitative methods (red - z-lib Compression; blue - BDM and orange - Entropy), it is shown (left) that they are highly insensitive to signal and highly sensitive to noise (16.5% of pixels randomly flipped). In that case, the correct partition size ($n=23$, so $n-9=14$) is modestly highlighted. However, when growing the original image (right) by a factor of 6 on each dimension (i.e. 1 pixel becomes a $6\times 6$ array), the methods become less sensitive to noise and more sensitive to signal amplification, with BDM significantly outperforming Compression, and Shannon Entropy showing sensitivity at up to 60% pixels flipped (hence about 30% of the original image) versus Compression and Shannon Entropy, that are about 50% sensitive. Downward spikes (right) are shown at $23\times 6-2=136$.
Investigating this further, Fig. 6 shows six bidimensional images along with their original numerical dimension. The values of the three complexity measures over possible partitions are shown below the original numerical dimension for ease of comparison. For presentation purposes, the values of BDM and z-lib b64encoded string have been divided by 1000 and then all the three metrics are shown in a logarithmic scale.
@@
The method is invariant with regard to linear transformations such as encoding, as shown in images 2 and 4 of Fig. 6. Drops in complexity (downward spikes) indicate candidate dimensions for the original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. BDM outperforms compression and entropy and correctly identifies the original message (image), encoding for 4 out of 6 images.
196                                   82                                     215
276                                 360                                 220
Figure 6: Six 2D images (labelled 1 through 6 going from left to right, top to bottom) very different in nature, including a demonstration of linear-transformation invariance conforming to the underlying theory (in this case, rotation of a mathematical formula). Size invariance has actually shown amplification of signal-to-noise difference. Under each image is its correct numerical (first) dimension. The values of the three complexity indexes over possible partitions are shown below the original numerical dimension for ease of comparison. Downward spikes indicate candidates for possible original partitions. In all cases, the correct dimension value is among the top three candidates, with BDM outperforming at indicating the top candidate in 4 out of 6 cases.
@@
Some of these image decodings for the space shuttle image are shown in Fig. 7. Each panel shows random realignments of the correct dimensions along with the original image (i.e., object) embedded into these new partitions, with their respective BDM values (without any scaling) displayed above each image. When compared to the other bidimensional spaces, the one that is most similar to the original message will have a lower complexity value.
Figure 7: Sequence of lowest complexity (BDM value on top in bits) partitions of a black and white image of a space shuttle, with the top right partition as the correct reconstruction and other candidate cases in which an alignment occurs at some multiple of the correct dimension.